40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1701     Entry time: Thu Jun 25 10:28:58 2009
Author: steve 
Type: HowTo 
Category: VAC 
Subject: Cryopump is regenerated 
The Cryopump's VC1 valve to IFO was closed yesterday.
The compressor  Helium pressure was 235 PSI. The cold head temp on H2 vapor pressure gauge was reading ~14 Kelvin,
The compressor and piston driver were turned off to let cold head warm up to room temp.
The flow path from Cryo to TP3 were checked to insure that only VC2 and V5 would be open for pumping.
VC2 valve was opened to TP3 through V5
Now as the Cryo was warming up while TP3 drag turbo pump was pumping away the accumulated ice, that was melting and vaporizing.
This is shown on one day the plot below.
To check outgassing rate of the Cryo pump after one day of pumping V5 was closed for 20 minutes.
The accumulation was 1.3 mTorr in 20 min
This means the Cryo is clean, it is ready to be started up in the future.
VC2 was closed to seal this condition.
The flow path between VC2, VM3, V7, V6 , VA6 and manual needle valve  would be pumped for one day through V5 to TP3 to clean up 
Attachment 1: Creg.jpg  56 kB  | Show | Show all
ELOG V3.1.3-